# Fish Tracking, Counting, and Behaviour Analysis in Digital Aquaculture: A Comprehensive Review

### v1

@@ II-B Fish tracking based on 3-dimensional visual information
3-D tracking methods offer advantages over 2-D tracking algorithms, as they can be used to study the behaviour of social animals and effectively address most occlusion problems. However, 3-D tracking also presents significant challenges due to the large number of fish, similar individual appearance, occlusion, and uncertainty of stereo matching.
@@
Figure 3: Three methods to measure the 3-D position of a fish in an aquarium.
@@
Two main types of 3-D tracking methods have been developed: ”shadow” and ”stereo” methods (as shown in Fig. 3). The ”shadow” method, which requires only one camera, uses the shadow of the fish projected onto the substrate as a second view of the shoal. By calculating the 2-D positions of the fish and its shadow, the 3-D position of the fish can be obtained through triangulation. However, this method becomes increasingly difficult as the number of fish increases and shadows may be obscured, as it requires detecting each fish and its corresponding shadow.
@@
Stereoscopic methods use multiple cameras to capture simultaneous images at different angles or a camera and a mirror [22]. Some researchers have developed platforms that use a single camera and mirror to obtain 3-D coordinates and automatically track fish [65, 66]. These methods calculate the centre coordinates of fish and combine the association of mirror view and direct view for tracking, addressing the problem of target loss caused by occlusion. However, they require high-precision equipment and may suffer from correspondence deviations due to the pixel centres of real and virtual fish not being at the same point. Moreover, these methods are not suitable for actual production environments.
@@
In theory, two cameras are sufficient for stereo imaging. 3-D tracking with two cameras involves obtaining the 2-D motion trajectory from the top view (larger viewing angle) and then performing 3-D matching of the top view tracking results with the feature points in the side view to obtain the object’s movement in 3-D space (as shown in Fig. 4) [67]. Three or more high-speed cameras are usually required to capture synchronous videos to track many objects, solve ambiguities, and avoid errors between objects. A study by [68] determined the location of fisheye under the top and side views using mixed Gaussian and Gabor models, respectively, and then obtained the 3-D motion trajectories of the objects by associating the top-view tracking results with the trajectories of two side views [40]. However, the detection effect has poor performance due to the difficulty in distinguishing the eye area characteristics of fish. Furthermore, analyzing fish movement behaviour in three views leads to complex equipment installation and reduces the accuracy of association and stereo matching [67].
@@
Figure 4: Three-dimensional trajectory of multiple fish in water tank via multi video tracking.
Occlusion remains one of the main challenges in 3D fish tracking, as it is in other MOT (Multiple Object Tracking) tasks. However, the frequency of occlusion has not been adequately measured in the current literature, with the complexity indicator of the datasets used in existing studies typically being the number of fish rather than an assessment of fish occlusion events. For instance, a demo video in [27] shows only 4 occlusion events within 15 seconds for a group of 10 fish.
@@
Current system evaluations assess parameters such as ID swaps, fragments, precision, and recall for the generated 2-D and 3D tracks without describing how these indicators are calculated. The lack of uniform indicators makes it difficult to fairly compare the methods presented in various studies. Furthermore, most of the literature does not provide open-source code and annotated data, limiting the repeatability of the results. A recent study by [69, 38] introduced a standard MOT evaluation framework for fish tracking, providing a good model for multi-target fish tracking. A unified evaluation standard should be introduced to ensure the fairness of fish multi-target tracking comparisons and facilitate progress in this field.
@@ II-C Fish tracking based on acoustic tag system
The Acoustic Tag System (ATS), a passive acoustic method of acoustic monitoring technology, has become an important means of monitoring fish trajectories and studying fish behaviour [70]. Unlike vision-based tracking methods, which rely on clear water conditions and sufficient lighting, ATS can provide reliable tracking data in challenging underwater environments, such as turbid waters or low-light conditions [71]. The appropriate acoustic tag (also called an acoustic signal transmitter) type and parameters are selected according to the size of the fish and the research period (as shown in Fig. 5) [72]. The application of the acoustic tagging system mainly includes the abundance assessment of fish resources, the swimming pattern of fish, the evaluation of habitat characteristics, the spawning site of fish, the survival situation of fish, and the behaviour differences of fish [73, 74]. However, acoustic tag monitoring technology is rarely used in aquaculture and has broad application prospects.
Using acoustic tag monitoring technology to monitor fish movement and behaviour trajectories, obtain real-time three-dimensional movement trajectory coordinates of the fish, and perform related data analysis and application is an advanced technique [75]. Compared with vision-based monitoring technologies, acoustic tag monitoring technology has the advantages of in-situ observation and simple data processing methods. However, this technology determines the location of the fish by receiving the sound wave signal sent by the acoustic tag on the fish, and the fish may die during the data monitoring process [76, 77]. Therefore, technicians must conduct real-time monitoring and data processing and analysis of monitoring data promptly to ensure the continuity and accuracy of the data.
Acoustic tag monitoring technology and vision-based tracking methods have their unique strengths and limitations. Vision-based methods can provide detailed information about fish appearance, shape, and motion, but they are limited by water clarity and lighting conditions. In contrast, acoustic tag monitoring technology can provide reliable tracking data in challenging underwater environments, but it lacks detailed visual information about fish appearance and behaviour. Combining these two modalities can help overcome their limitations and provide a more comprehensive understanding of fish behaviour and movement patterns.
Figure 5: Fish tracking method based on acoustic tag[78].
@@ II-D Tracking evaluation metrics
Multi-target tracking evaluation indices directly reflect an algorithm’s tracking ability, and the MOTchallenge official multi-objective tracking evaluation indicators [79] provide a standardized framework for assessment. Key metrics include Multiple Object Tracking Accuracy (MOTA) and Multiple Object Tracking Precision (MOTP).
The MOTA combines three sources of errors to evaluate a tracker’s performance, follow as:
@@
where $FN_{t},FP_{t},ID_{sW_{t}},GT_{t}$ represent the number of false negatives, false positives, identity switches, and ground truth targets in frame $t$ correspondingly.
@@
To better capture the specific challenges of tracking fish populations, some literature has introduced additional metrics, such as Correct Tracking Ratio (CTR) and Correct Identification Ratio (CIR). CTR measures the percentage of correctly tracked frames for individual fish.
@@
$CIR$ represents the probability of correctly identifying all fish after an occlusion event:
@@
In addition to those metrics, tracking speed is another important factor to consider when evaluating fish-tracking algorithms, especially for real-time applications. Some common metrics for measuring tracking speed include frames per second (FPS) and processing time per frame. FPS indicates the number of frames a tracking algorithm can process in one second while processing time per frame measures the average time taken to process a single frame. Higher FPS and lower processing time per frame are desirable for efficient and real-time tracking performance. These metrics offer a valuable foundation for evaluating fish tracking performance, comprehensively assessing various errors, fish-specific challenges, and tracking speed. However, they may not always capture the full complexity of fish-tracking scenarios and can be limited by the lack of widespread adoption and the need for detailed ground-truth annotations.
To drive advances in this field, researchers should work towards developing more specialized metrics and evaluation protocols that consider the specific requirements and challenges of fish tracking applications. By combining these metrics with careful consideration of the diverse underwater environments in which fish tracking algorithms must operate, researchers can work towards more comprehensive and standardized evaluation practices that fully characterize the robustness, generalizability, and efficiency of these algorithms.
@@ III-A Fish counting methods based on sensor technology
Sensor-based counting devices are usually divided into resistance counters and infrared counters. Infrared counters detect infrared signals, which are electromagnetic waves with wavelengths between 760 nm and 1 mm [80]. Counting based on infrared counters requires a tunnel structure to limit the movement of the fish. When a fish passes between the infrared transmitter and the receiver, the counting is completed [81, 82]. Although infrared sensors can count in smaller areas of space, their performance is affected by water depth and turbidity. At a depth of 17.9 centimetres (cm) in pure water, the intensity of the infrared light drops to 50% [83], and the presence of suspended particles can further degrade the performance of infrared counting at high turbidity levels [84]. In addition to environmental factors, the accuracy of infrared sensor counting devices is susceptible to the pass rate of fish, often resulting in an underestimation of the number of fish [85, 86]. This may be due to the slow swimming of some fish, confusion when two or more fish enter the scanner unit simultaneously, and the reluctance of some fish to leave the device after entering the light tunnel, resulting in repeated scanning. Despite these limitations, infrared light can work in the dark, and the accuracy of counting can be improved by subsequent software algorithms, such as multiple object tracking (MOT) algorithms, which can solve false counting from multiple targets [84, 87, 88].
Resistivity counters, another type of sensor-based counting method, work by detecting changes in resistance when a fish passes between two electrodes [89, 90]. Like infrared sensor counting devices, electronic resistivity counters require the fish to pass through a specific tunnel and have similar disadvantages, such as repeat counts when a fish swims multiple times in the channel and missing counts when the number of fish is large [91]. However, electronic resistivity counters are suitable for limited lighting and long, narrow river channels while detecting non-destructively and without requiring specific lighting conditions [92].
Although both infrared and resistivity fish counters have limitations and may underestimate fish pass rates, they offer valuable tools for non-invasive fish counting in various environments. Future research should focus on developing and improving these technologies to enhance their accuracy and reliability. Potential avenues for improvement include modifying resistivity counters and exploring alternative sensors [93]. By addressing the current challenges and refining these sensor-based counting methods, researchers can provide valuable tools for effective fishery management and conservation efforts.
III-B Counting methods based on computer vision technology

### v2

@@ II-B Fish tracking based on 3-dimensional visual information
3-D tracking methods offer advantages over 2-D tracking algorithms, as they can be used to study the behaviour of social animals and effectively address occlusion problems. However, 3-D tracking also presents significant challenges due to the large number of fish, similar individual appearance, occlusion, and uncertainty of stereo matching.
@@
Figure 3: Three methods to measure the 3-D position of a fish in an aquarium (This figure was adapted from [11]).
@@
Two main types of 3-D tracking methods have been developed: “shadow” and “stereo” methods (as shown in Fig. 3). The “shadow” method, which requires only one camera, uses the shadow of the fish projected onto the substrate as a second view of the shoal. By calculating the 2-D positions of the fish and its shadow, the 3-D position of the fish can be obtained through triangulation. However, this method becomes increasingly difficult as the number of fish increases and shadows may be obscured, as it requires detecting each fish and its corresponding shadow [83].
@@
Stereoscopic methods use multiple cameras, a camera and a mirror to capture simultaneous images at different angles [11]. Some researchers have developed platforms that use a single camera and mirror to obtain 3-D coordinates of fish [84, 85]. These methods calculate the centre coordinates of fish and combine the association of mirror view and direct view for tracking, addressing the problem of target loss caused by occlusion. However, they require high-precision equipment and may suffer from correspondence deviations due to the pixel centres of real and virtual fish not being at the same point. Moreover, these methods are not suitable for actual production environments.
@@
In theory, two cameras are sufficient for stereo imaging. 3-D tracking with two cameras involves obtaining the 2-D motion trajectory from the top view with a larger viewing angle and then performing 3-D matching of the top view tracking results with the feature points in the side view to obtain movement of the object in 3-D space (as shown in Fig. 4) [86]. To track many objects, three or more high-speed cameras usually are needed to capture synchronous videos to mitigate the ambiguities and errors in distinguishing the objects. For example, Wang et al. [87] determined the location of fisheye under the top and side views using mixed Gaussian and Gabor models, respectively, and then obtained their 3-D motion trajectories by associating the top-view tracking results with the trajectories of two side views [39]. However, this method is limited due to the difficulty in distinguishing the characteristics from the eye area of fish. Furthermore, analyzing fish movement behaviour in three views requires the installation of complex equipment and the association and stereo matching between the views [88].
@@
Figure 4: Three-dimensional trajectory of multiple fish in water tank via multi video tracking (This figure was inspired by [86] with minor changes).
@@
Occlusion remains one of the main challenges in 3D fish tracking, as it is in other Multiple Object Tracking (MOT) tasks. However, the frequency of occlusion has not been adequately measured in the current literature, with the complexity indicator of the datasets used in existing studies typically being the number of fish rather than an assessment of fish occlusion events. For instance, a demo video in [27] shows only 4 occlusion events within 15 seconds for a group of 10 fish. Current system evaluations assess parameters such as ID swaps, fragments, precision, and recall for the generated 2-D and 3D tracks without describing how these indicators are calculated. The lack of uniform indicators makes it difficult to fairly compare the methods presented in various studies. Furthermore, most of the literature does not provide open-source code and annotated data, limiting the reproducibility of the results. A recent study by [89, 37] introduced a standard MOT evaluation framework for fish tracking, providing a good model for multi-target fish tracking. A unified evaluation standard should be introduced to ensure the fairness of fish tracking comparisons and facilitate progress in this field.
Figure 5: Fish tracking method based on an acoustic tag (This figure was reproduced from[90] with minor changes).
@@ II-C Fish tracking based on acoustic tag system
The Acoustic Tag System (ATS), a passive acoustic monitoring technology, has become an important means of monitoring fish trajectories and studying fish behaviour [91]. Unlike vision-based tracking methods, which rely on clear water conditions and sufficient lighting, ATS can provide reliable tracking data in challenging underwater environments, such as turbid waters or low-light conditions [92]. The appropriate type and parameters of the acoustic tag (also called an acoustic signal transmitter) are selected according to the size of the fish and the research period (as shown in Fig. 5) [93]. The applications of ATS are diverse, including fish resource abundance assessment, swimming pattern analysis, habitat evaluation, spawning site identification, survival rate estimation, and behavioural differentiation [94, 95]. Kolarevic et al. [96] demonstrated the use of acoustic acceleration transmitter tags for monitoring Atlantic salmon swimming activity in recirculating aquaculture systems (RAS), while Fore et al. [97] found it feasible to monitor fish depth in real-time in commercial fish farms using acoustic telemetry.
Acoustic tag systems have been used to monitor fish movement and trajectories, obtain the three-dimensional coordinates of the fish in real-time, and perform related data analysis and application. A notable example of an ATS application involves 682 subyearling Chinook salmon tagged with injectable acoustic transmitters in the Snake River, Washington State [90]. This study generated over 5 million clean detections and 403,900 3-D trajectories, helping to assign fish to passage routes through a dam and expanding understanding of near-dam fish behaviour [90]. Leclercq et al. [98] and Munoz et al. [99] evaluated the feasibility of using passive acoustic telemetry to monitor the welfare of fish in sea cages on an industrial scale, and the experimental results showed that the acoustic tracking system proved to be an effective tool for fish behaviour analysis. However, the application of acoustic tags in commercial aquaculture faces significant challenges. Macaulay et al. [100] systematically reviewed studies using tags to monitor farmed fish behaviour and found concerning results regarding tag-associated mortality. Mortality was 10 times higher in marine caged fish compared to tanks. Moreover, mortality in marine caged fish increased significantly in longer trials, ranging from 4% in single-day trials to 36% after 100 days.
Despite these challenges, ATS offers unique advantages in monitoring fish movement and obtaining real-time three-dimensional trajectory coordinates [101]. It excels in in-situ observation and simplifies data processing over vision-based methods. However, the technology’s reliance on tags attached to fish can potentially impact on fish health and survival [102, 103], necessitating careful monitoring and prompt data analysis. Future research should focus on improving tag design, developing standardized reporting protocols, investigating mortality causes in marine caged fish, and exploring less invasive alternatives. By addressing these challenges, ATS could become a more viable tool for understanding and improving fish welfare and productivity in aquaculture, though its use should be limited currently and results interpreted cautiously in commercial settings.
@@ II-D Tracking evaluation metrics
Multi-target tracking evaluation indices directly reflect an algorithm’s tracking ability. The MOT Challenge official multi-objective tracking evaluation indicators [104] provide a standardized framework for performance assessment. Key metrics include Multiple Object Tracking Accuracy (MOTA) and Multiple Object Tracking Precision (MOTP).
The MOTA combines three sources of errors to evaluate a tracker’s performance, as follows:
@@
where $FN_{t},FP_{t},ID_{sW_{t}}$, and $GT_{t}$ represent the number of false negatives, false positives, identity switches, and ground truth targets in frame $t$, respectively.
@@
To better capture the specific challenges of tracking fish populations, some literature has introduced additional metrics, such as Correct Tracking Ratio (CTR) and Correct Identification Ratio (CIR). CTR measures the percentage of correctly tracked frames for individual fish as follows,
@@
while $CIR$ represents the probability of correctly identifying all fish after an occlusion event:
@@
In addition to those metrics, tracking speed is another important factor to consider when evaluating fish-tracking algorithms, especially for real-time applications. Some common metrics for measuring tracking speed include frames per second (FPS) and processing time per frame. FPS indicates the number of frames a tracking algorithm can process in one second while processing time per frame measures the average time taken to process a single frame. Higher FPS and lower processing time per frame are desirable for efficient and real-time tracking performance. These metrics offer a valuable foundation for evaluating fish tracking performance, comprehensively assessing various errors, fish-specific challenges, and tracking speed. However, they may not always capture the full complexity of fish-tracking scenarios and can be limited by the lack of widespread adoption and the need for detailed ground-truth annotations. There is a potential scope to develop more specialized metrics and evaluation protocols by considering the specific requirements and challenges of fish tracking applications.
TABLE II: Comparison of fish tracking methods
Method Advantages Disadvantages Welfare Implications Suitable Environments
2D Visual
[37, 105]
[78, 49]
Non-invasive
Cost-effective
Real-time tracking
Limited to surface or shallow waters
Affected by water turbidity
Occlusion issues in dense populations
Minimal stress
No physical interaction
Clear water ponds
Shallow tanks
Surface monitoring in sea cages
3D Visual
[89, 106]
[107, 108]
Accurate spatial tracking
Works in deeper waters
Better for dense populations
Higher computational requirements
More complex setup
Costly equipment
Minimal stress
Potential light disturbance
Deep tanks
Sea cages
Research facilities
Acoustic Tags
[109, 110]
[111, 112]
Works in turbid waters
Long-range tracking
Effective in open waters
Invasive (if using tags)
Potential signal interference
Higher initial cost
Stress from tagging procedure
Potential behavioral changes
Open sea farms
Large lakes
Deep water environments
II-E Comparative analysis of tracking methods
2D visual methods are ideal for controlled environments with clear water, offering a balance between cost-effectiveness and non-invasiveness. However, they struggle in turbid or deep waters and occlusions in dense populations. 3D visual tracking provides more comprehensive data but requires more sophisticated equipment and setup, making it more suitable for research-intensive or high-value production environments. Acoustic tag methods excel in challenging water conditions but involve more invasive procedures that may impact fish welfare (as shown in Table II). The choice of a method often depends on the specific aquaculture environment. For example, shallow freshwater ponds might benefit more from 2D visual systems, while deep-sea cages may require acoustic or advanced 3D visual setups.
Regional preferences for tracking methods vary significantly. In Asia, 2D and 3D visual tracking is used predominantly, especially in countries like China and Japan [113, 114]. This preference is driven by the prevalence of intensive ponds, tank aquaculture, and strong technological infrastructure. A balanced approach is used in Europe, with significant adoption of all three methods. In North America, the integration of advanced AI with tracking systems is often employed, particularly by combining visual and acoustic tag methods for comprehensive monitoring [93]. In Oceania, acoustic tracking is favoured due to the prevalence of open-water aquaculture, especially in tuna farming [115]. Environmental conditions, regulatory frameworks, economic factors, and available technological infrastructure shape the trends for adoption. Developing regions in Africa and parts of Asia often lack access to advanced tracking technologies, highlighting the need for cost-effective solutions for small-scale farmers.
Animal welfare is an increasingly important consideration for the selection and development of tracking methods [116]. Visual methods generally cause minimal direct stress but may alter behaviour due to changes in lighting or the presence of equipment [117]. Acoustic tagging can cause initial stress during the tagging procedure and may affect long-term behaviour due to the presence of the tag [110, 112]. All methods can potentially cause stress if they require changes to the fish’s environment or handling of the fish. Research on less invasive techniques includes the development of smaller, lighter acoustic tags to minimize the impact on fish movement and behaviour [3], advancements in computer vision and AI to improve tracking accuracy without the need for physical tags [118], and exploration of environmental Deoxyribonucleic Acid (DNA) or eDNA (Genetic material collected from environmental samples) methods for non-invasive population monitoring [119]. Balancing data collection needs with fish welfare involves implementing rotation systems in tagging studies to reduce the duration of tag attachment for individual fish, using hybrid methods to reduce the reliance on any single, potentially stressful technique, and developing guidelines for the ethical use of tracking technologies in aquaculture, considering factors such as tag size relative to fish size, duration of studies, and recovery periods [120, 119]. The aquaculture industry is trending towards methods that minimize stress and interference with natural behaviours while providing effective management data. This shift is driven by both ethical concerns and the recognition that improved welfare contributes to better product quality and farm productivity.
II-F Real-world applications of fish tracking
Real-world applications of fish tracking primarily span two domains: commercial aquaculture monitoring in controlled tank environments and marine ecological research in natural underwater settings. While aquaculture applications focus on fish behaviour and health monitoring through overhead cameras, underwater tracking faces more complex challenges including varying visibility, coral occlusions, and unpredictable fish movements. A recent study by Zhao et al. [69] demonstrates the potential of advanced tracking and behaviour analysis techniques in aquaculture settings. The researchers developed a method combining an improved ByteTrack algorithm “Fish amendments (FA)-ByteTrack” with a spatiotemporal graph convolutional network (ST-GCN) to assess fish appetite in a controlled environment. In their experiment, they used a circular tank (1.5 m × 1.5 m × 1 m) with 120 large-mouth bass, each weighing 400-500 grams. This setup aimed to simulate the complex conditions of real aquaculture environments, including fish overlapping and occlusions. The system used a high-resolution camera (1920x1080 pixels) positioned 1.6 m above the water surface, recording at 60 frames per second. This method not only effectively solves the problems of bubble masking, intraclass variation and cross-occlusion but also achieves high-precision appetite assessment. Compared with that of ByteTrack, the average Multiple Object Tracking Accuracy (MOTA) of FA-ByteTrack increased by 21.3%. Moreover, the tool achieved 98.47% accuracy in appetite assessment (as shown in Fig. 6 ).
Figure 6: Comparison of ByteTrack and FA-Bytetrack in controlled environment (This figure was reproduced from[69]).
However, it is worth noting that while this study represents a significant advancement, it was conducted in a controlled environment that does not fully reflect the challenges of real-world aquaculture settings. In commercial aquaculture operations, factors such as larger tank sizes, higher fish densities, varying water conditions, and different species could significantly impact the performance of such systems. Moreover, the study relied on a pre-trained YOLOv8 detector, which may need a large amount of high-quality labelled data for training detectors. Labelling them is time-consuming and costly.
The FISHTRAC dataset [105] exemplifies the challenges of real-world fish tracking in underwater environments, as shown in Fig. 7. Unlike controlled aquarium settings, natural underwater environments present unique challenges. Fish exhibit unpredictable movements, and rapidly changing appearances, and are frequently occluded by coral or other marine life. Moreover, fish often actively attempt to evade or hide from divers, adding another layer of complexity to the tracking task. A recent work by Mandel et al. [105] proposed a detection confidence-driven multi-object tracking to recover reliable tracks from unreliable detections. The study relied on a limited dataset of only 1800 images labelled as “fish” from the Google Open Images Dataset [121], many of which were not from real underwater environments. The limited scale of this dataset resulted in a detector with mediocre performance, highlighting the real-world challenges of applying deep learning methods to specialized domains like underwater fish tracking.
Despite these limitations, the Robust Confidence Tracking (RCT) algorithm [105] showed promising results. The key strength of RCT lies in its ability to fuse low-confidence detections with a motion model and single object tracker, maintaining consistent tracks even when high-confidence detections are sparse. This capability is especially valuable in challenging underwater conditions where high-quality detections are often unavailable. The study highlights the potential of RCT for practical applications in marine biology research and ecological monitoring, particularly in scenarios with scarce labelled data and challenging detection conditions. Advanced tracking algorithms based on Bayesian filters [122], such as particle filters [123] and Bernoulli filters [124], offer additional tools for trajectory smoothing and handling outliers or mis-detection. These methods, widely used in audio-visual speaker tracking, offer the potential to be applied in underwater environments. Future research directions include developing online versions of these algorithms for edge computing, improving performance in high-density scenarios through Bayesian incorporation of appearance information, and exploring adaptive approaches that adjust tracker behaviour based on detection quality. These advancements aim to bridge the gap between controlled studies and the complex realities of commercial aquaculture, paving the way for more robust and practical fish tracking and behaviour analysis systems.
Figure 7: FISHTRAC dataset in underwater environments (This figure was reproduced from [105]).
@@ III-A Fish counting methods based on sensor technology
Sensor-based counting devices are usually divided into resistance counters and infrared counters. Infrared counters detect infrared signals, which are electromagnetic waves with wavelengths between 760 nm and 1 mm [125]. Counting based on infrared counters requires a tunnel structure to limit the movement of the fish. When a fish passes between the infrared transmitter and the receiver, the counting is completed [126, 127]. Although infrared sensors can count in small areas of space, their performance is affected by water depth and turbidity. At a depth of 17.9 centimetres (cm) in pure water, the intensity of the infrared light drops to 50% [128], and the presence of suspended particles can further degrade the performance of infrared counting at high turbidity levels [129]. In addition to environmental factors, the accuracy of infrared sensor counting devices is susceptible to the pass rate of fish, often resulting in an underestimation of the number of fish [130, 131]. This may be due to the slow swimming of some fish, the confusion when two or more fish enter the scanner unit simultaneously, and the reluctance of some fish to leave the device after entering the light tunnel, resulting in repeated scanning. Despite these limitations, infrared light can work in the dark, and the accuracy of counting can be improved by subsequent software algorithms, such as multiple object tracking (MOT) algorithms, which can solve false counting from multiple targets [129, 132, 133].
Resistivity counters, another type of sensor-based counting method, work by detecting changes in resistance when a fish passes between two electrodes [134, 135]. Like infrared sensor counting devices, electronic resistivity counters require the fish to pass through a specific tunnel and have similar disadvantages, such as repeating counts when a fish swims multiple times in the channel and missing counts when the number of fish is large [136]. However, electronic resistivity counters are suitable for limited lighting and long, narrow river channels while detecting non-destructively and without requiring specific lighting conditions [137].
Although both infrared and resistivity fish counters have limitations and may underestimate fish pass rates, they offer valuable tools for non-invasive fish counting in various environments. Future research should focus on developing and improving these technologies to enhance their accuracy and reliability. Potential avenues for improvement include modifying resistivity counters and exploring alternative sensors [138]. Researchers can provide valuable tools for effective fishery management and conservation efforts by addressing the current challenges and refining these sensor-based counting methods.
III-B Fish counting methods based on computer vision technology